OpenStudy (kerrionlo):

What is the solution to the equation 1/2x+3=2/3x+1?

OpenStudy (kevin):

step 1: remove denominator by multiply both side with denominator least common multiple step 2 : move number contains x to left and the other to right step 3 : simplify the equation done

OpenStudy (cmdrwinters):

To solve this, we must first isolate the variable on one side of the equation. \[\frac{ 1 }{ 2 }x + 3 = \frac{ 2 }{ 3 }x + 1\] First step in this process would be to combine the like terms on their own side of the equation. \[\frac{ 1 }{ 2 }x + 3 - 3= \frac{ 2 }{ 3 }x + 1-3 \] \[\frac{ 1 }{ 2 }x = \frac{ 2 }{ 3 }x -2 \] \[\frac{ 1 }{ 2 }x -\frac{ 2 }{ 3 }x= \frac{ 2 }{ 3 }x - \frac{ 2 }{ 3 }x -2 \] At this point we can find a common denominator to make simplifying the fractions much easier. The lowest common denominator for these fractions would be 6, so we can make it \[\frac{ 3 }{ 6 }x -\frac{ 4 }{ 6 }x= \frac{ 4 }{ 6 }x - \frac{ 4 }{ 6 }x -2 \] \[ -\frac{ 1 }{ 6 }x= -2 \] Now, we just divide both sides by the fraction to find the variable. \[ -\frac{ 1 }{ 6 }x / -\frac{ 1 }{ 6 } = -2/-\frac{ 1 }{ 6 } \] \[ x =12\] Which is our answer. Hope that helped! :)

jhonyy9 (jhonyy9):

please use parentheses bc. not is clearly what mean denominator 2x or 2x +3 and on the other side 3x or 3x +1 ?